There are 13 animals in the barn. Some are chickens and some are pigs. There are 40 legs in all. How many of each animal are there?
Found 2 solutions by ikleyn, ewatrrr:
Answer by ikleyn(52867)

.

a TWIN problem was solved for you 4 minutes ago at this link

https://www.algebra.com/algebra/homework/Systems-of-equations/Systems-of-equations.faq.question.1174011.html

https://www.algebra.com/algebra/homework/Systems-of-equations/Systems-of-equations.faq.question.1174011.html


Consider it as your TEMPLATE.


Having this template in front of you, solve the current problem in the same way.


E n j o y (!)


--------------

To see many other similar solved problems, look into the lessons
    - Problem on two-wheel and three-wheel bicycles
    - Problem on animals at a farm
    - Problem on pills in containers
    - What type of problems are these?
in this site.




Answer by ewatrrr(24785)

 
Hi,
There are 13 animals in the barn. Some are chickens and some are pigs. 
There are 40 legs in all. How many of each animal are there?
 Let x be the number of pigs and (13-x) the number of chickens
  4x + 2(13-x) = 40 legs
  4x + 26 - 2x = 40
        2x = 14 
          x = 7  is the number of pigs 
              6  is the number of chickens 13-7
checking our work:
    4*7 + 2*6= 28 + 12 = 40      
Wish You the Best in your Studies.




/\/\/\/\/\/\/\/

To tutor @ewatrrr

    Please look into the last line in your post source code.


    It's me who added this line to your code.


    May I ask you PLEASE to provide this line (such line) at the end of each your post ?


    It serves to CLOSE  the font options  of your source code.


    Otherwise, these options extend their influence to the posts of other tutors on the same subject.


    THANK YOU.


    @ikleyn


It is, probably, my attempt #7 to bring this message to your attention.

I placed it in my posts several times today, but it seems to me that you do not see anything here at the forum.
@ikleyn, am aware of your concern, did not know how to address it manually.
Using what You now provided, leaves the spacing of lines too close.
Prefer more spacing. Perhaps you might have another suggestion.

...............

Sure.

Simply make one empty line between your text lines. That is all.